OpenStudy (anonymous):

Solve by giving square roots: 1) -2v^2=28 2) -10x^2=90

OpenStudy (anonymous):

(1) i sqrt14

OpenStudy (anonymous):

could you explain the process, please?> I wish to learn

OpenStudy (anonymous):

(1) -2v^2=28 --- given v^2=-14 --- multiplicative inverse property v =sqrt -14 --- square root property v = sqrt(14)(-1) given that sqrt of -1 is i (imaginary number) v = i sqrt 14 ; since 14 has no perfect square factor

OpenStudy (anonymous):

and the second one,? Very well explained thank you btw

OpenStudy (anonymous):

(2) -10x^2=90 x^2=-9 (multiplicative inverse property) x =sqrt(9)(-1) (factor the radicand) ;9 is perfect square x = +-3i

OpenStudy (anonymous):

Great thanks !

OpenStudy (anonymous):

glad to help. .